Evolution and Jung’s Theory of the Unconscious

Anthony K. Anderson

May 2007 (revised Dec 2008)

My soul gives me these forms to sing, transformed To bodies strange and new: O gods inspire These new beginnings, for you have also changed [.]

— Ovid (Metamorphoses, lines 1–3)

The wisdom of pre-modern philosophy and the pre-modern psyche is now beyond the reach of many people, not because of any imprecision on the part of our philologists and archeologists, but because our minds cannot easily make room for those old ideas. The theory of evolution now informs our understanding of everything in nature, including ourselves, and by doing so it seems to undermine any alternative point of view which is based on an incompatible view of our origin. In place of a divine origin we now seem to have a natural, some would say pathological, origin. Maybe it would be better for us to take the high path that William James proposes for the philosophy and psychology of religion. “The plain truth,” he writes, “is that to interpret religion one must in the end look at the immediate content of the religious consciousness.”^{1} However, the religious consciousness of every age and every culture is like a different puzzle, though perhaps cut from the same original image, and we cannot easily use old pieces to complete our own picture of the world: the pieces we have are different now.

In the psychology of Carl Jung, the “Unconscious” is one central piece of the modern puzzle. As a modern analogue of the traditional idea of the Divine, it provides psychology with a new way to understand religious experience. But the concept of the Unconscious is no less difficult to comprehend than its traditional counterparts. Even as it is applied in Jung’s clinical practice, it precipitates a variety of theoretical difficulties. And from a philosophical or religious point of view, the problems are also great. What is the Unconscious? What does it do? What is its origin? How could it have acquired the wisdom it seems to have? How much does it actually resemble the traditional conception of the Divine? As an author Jung always emphasizes his “psychological empiricism” and his disregard for both biological and metaphysical causal explanations.^{2} Jung would probably wish to say simply that the Unconscious is wise if it acts wise, regardless of whether we can explain the source of its wisdom by means of some hypothesis about its origin. His theory of psychology certainly seems to be quite compatible with a biological explanation of the Unconscious, and of religious experience in general, but nevertheless Jung is always reluctant to stand firmly in this reduction. He even seems to believe that some religious experiences happen before, above, or behind our conventional experiences of the world in which the categories of

1William James, The Varieties of Religious Experience (in Writings: 1902-1910), 20 footnote.

2C.G. Jung, Memories, Dreams, Reflections, 349-353. In his psychological empiricism Jung follows the example of William James, cf. The Varieties of Religious Experience, 25

Anderson 2

mind/body, material/spiritual, etc, apply.^{3} Still, Jung makes consistent use of biological analogies when he talks about the archetypes in the Unconscious, comparing them directly to the intelligent but unconscious patterns of behavior we see in animals.^{4} So I think it does make sense to look at just how much of his theory of the Unconscious can be accounted for within the scope of orthodox evolutionary theory. Briefly, it is the aim of this essay to answer the question: How did the Unconscious become wise? It will be helpful first to consider design processes in general and evolution specifically. After that we will take up the question of the Unconscious, in light of its biological origin.

Preliminary Discussion of Design Processes

The recent controversy over the theory of “intelligent design” illustrates an important ambiguity in the way we talk about design in nature. It seems as though the word “design,” by its definition, refers to something with intelligence, purpose, or some kind of wisdom in it or imprinted upon it. So when we speak of evolutionary design processes that are blind and unintelligent, it is important to be clear about the sense in which objects produced by these processes should qualify as “designs,” rather than just something surprising that happens. In fact, “something surprising that happens” is a very accurate description of evolution, and if it were possible for evolutionary theorists to avoid talking about “design” altogether, much confusion might be avoided. This semantic difficulty seems to be at the heart of the intelligent design debate, as well as much of the popular misunderstanding of evolutionary theory. But even if we were to choose a different word . “evolutionary production,” “development,” etc – rather than “design,” a significant ambiguity would still remain: evolutionary theorists are focusing on the aim and intentions of the designer when they label evolution blind and unintelligent, and they are focusing on their own perception of the end product of the process when they call it a design process. For a long time it had not occurred to people that apparent design could arise from an unintelligent process. Even today it is hard to understand how it happens, and what it means for such a thing to happen. For that reason, I think it would be helpful to stand back and make a more general analysis of what design processes consist of and how they work.

The primary difficulty in this analysis is that every design can be looked at in two ways. First, it is possible to examine the method that is used for deciding how the product will work, how it will be made, what it will be made out of, etc; secondly it is possible to examine the actual product itself, what its parts are, what it does. The methods used in the first stage often leave a mark on the actual product, so that it is sometimes possible to infer how the design was planned merely by examining the product; that however is not always so easy, and confusion arises as a result of these two viewpoints. One way to lessen this confusion is to think of design not as something that is inherent in objects, but rather as something that we humans see in objects as a result of the way we look at them. The recognition of design in nature is, in this sense, like the recognition of force in nature. A physicist infers the presence of a force when a physical body deviates from inertial motion. However, inertial motion is only a theoretical construction. According to Newtonian physics, for example, there is a gravitational force that causes an object to fall, contrary to its expected uniform rectilinear motion. But according to General Relativity, it is natural for an object

3For a general discussion see Jung’s essay “Synchronity: An Acausal Connecting Principle,” in Collected Works Vol.8

4C.f. C.G. Jung, “The Concept of the Collective Unconscious,” in Collected Works Vol. 9 Part 1, 43-44

Anderson 3

to “fall” when it is close to a massive object; because there is no deviation from inertial motion, there is no corresponding force. We can think of design in a similar way. So long as the origin of a particular object is partially unknown to us, so long as the object’s existence surprises us, we infer that the object has been designed. The concept of design, like that of force, is both intuitive and useful, and we can by no means do without it at present. But it is helpful to keep in mind, as we discuss design processes, that “design” is not something fixed, but rather a particular way of experiencing nature in its greater uniformity.

The design process that is most familiar to us, which could be called “human-like” design, I would categorize as having two primary properties: it is linear, and it has an aim. Consider a typical human-made machine. Each part in the machine is set into motion by one event, and results in one event. If you ask “What does this part do?” you will get an answer in the form “When A happens, this part does B.” It seems that humans tend to think most easily in terms of this kind of design. Each part is there because the inventor wanted something to happen at a specific time and put the part there for that purpose. Each part has a single purpose, or meaning, which is what I mean by linearity. It seems to me that the word “design” is most often used to mean this kind of linear functionality. In addition, the whole product is made in order to do something. This overarching motivation shapes the whole design process, and the design can be evaluated as a success or failure by this standard. This is what I mean by having an “aim.” Many designs do not fall strictly within this class. Most products have a variety of uses, and their parts do not function in a strictly linear fashion. For example, you can sit on a chair, but you also can use it in a philosophical demonstration. Here again it is helpful to think of design not as something inherent in the object, but rather as a way of looking at the object. Someone who thinks in terms of human-like design will see a chair as an object made for sitting on. The seat does one thing, and so do the back and legs. Most likely the designer himself thought of the chair just this way.

Darwinian evolution is more complicated, but I would venture to categorize it primarily as nonlinear production without aim. It is certainly true that some traits of biological organisms do seem to have one clear function. Wings are for flying, the heart is for pumping blood, pollen is for reproducing, etc. It makes some sense that for most of human history we have thought that nature has an anthropomorphic designer. For biologists however, it has started to make more sense to look at traits as having nonlinear aspects. The function of a particular trait is not limited to the linear vision of a particular designer at a particular time; in the course of the extensive time of evolutionary history, a trait may start to interact with any surrounding trait in many ways under many different conditions. It ceases to be possible to assign the trait a particular purpose or meaning. It might be possible to generate a long list of different functions of the same trait – different responses to different questions the environment can put to the organism – but even that might be misleading. There is never a point in the evolutionary process when the design needs to lay itself out on the table for inspection. Its development all happens on the run, and it is never used in the same way twice; in fact it is never used at all – at least not in the way we think of using human designs. Also, in contrast to human-like design, there seems to be no overarching motivation shaping this design process. Some would say that “survival of the fittest” or a more refined statement of the same idea should qualify as an overarching guiding principle. The question of whether this should qualify as “aim” is hardly unimportant; but there are at least two other permutations of the design process, linear design without aim, and nonlinear design with aim, which are clearly different from this one. Therefore I would prefer to maintain my somewhat ambiguous classification of Darwinian evolution as nonlinear design without aim.

Anderson 4

The case of linear design without aim does not seem particularly interesting. But the final case, nonlinear design with aim, provides a variety of useful examples. Because of the imprecision of my classification scheme, many of these seem to lie somewhere between true linear and nonlinear design. In the case of solutions to optimization problems in computer science, for example, evolutionary algorithms are sometimes used to find solutions more efficiently. The designer does not attempt to foresee how the algorithm will manage information at each step, but rather he allows the algorithm to gather whatever information it can. The result is that the algorithm has some of the flexibility characteristic of nonlinear design. Nevertheless, such an algorithm is still used in a linear fashion. If it is well designed, it takes a single input and reliably gives a correct output. There is also the (at present) theoretical problem of quantum computing, in which information is stored and managed in quantum superposition in a way that allows for great efficiency. In this case, each part of the physical machine literally does not have a single state (or meaning), but stores a variety of different messages in superposition in the same place. Quantum computers, though nonlinear in form, could still be used in a linear fashion; if well designed, they could be relied upon to give correct outputs with at least predictable frequency. The increasingly widespread study and use of nonlinear design with aim shows that nonlinear design processes in general – such as evolution – can serve not only as blind substitutes for linear design, but rather can at times far exceed them in terms of efficiency, adaptability, and general capacity.

Creationists and evolutionary theorists, despite their differences, are alike in that both of their theories of creation posit a non-human designer, and in both cases it seems to be hard for them to avoid taking an overly anthropomorphic view of these creators (though evolutionary theorists try harder to avoid it). We only know God and nature through our own human experiences, and so our knowledge of both is focused on humans. But it would probably be fair to say that both God and nature are in their essences, and even in many of their actions in the world, utterly alien and inhuman. In a way I think it is the overly anthropocentric view of creation as a whole that fuels the debate over intelligent design. Both creationists and evolutionists wish to emphasize how “mindless” and “ungodly” the orthodox evolutionary design process is, to show how dissimilar it is to a human design process. At the same time, they interpret designs in nature as if they can be understood in a linear way. This kind of thinking is exemplified in the debate over irreducible complexity. The most frequent question you hear debated is: “Is it possible that a blind nonlinear evolutionary process could have created this particular trait?” Most of the time, the more apt question regarding nature is: “Is it possible that a linear design process could have created this trait?”

Creationists seem to think that nature loses its meaning and purpose if we stop believing that God reached down out of the sky and molded the world with his five fingers. In a way they are right, but they are also missing out on the less anthropomorphic way of thinking about God and creation that evolutionary theory has provided them with. The great efficiency and adaptability, as well as the general style, of nonlinear design processes seem to have much in common with what theology has had to say about the intelligence of God. Really, it is the apparent aimlessness of evolutionary theory that seems to be the most radical departure from traditional views of God and creation, and the greatest challenge to religious morality, theories of natural law, and the like. It is at this point that Jung’s theory of the Unconscious becomes relevant, suggesting an interpretation of this difficulty and the possibility of partially reconciling traditional religious views with modern views of evolutionary design.

Anderson 5

Jung’s Theory of the Unconscious

The human psyche was traditionally thought to have three distinct parts: instinct, reflection, and spirituality. These three parts correspond roughly to the three traditional parts of the world: the material, the human, and the divine. Monotheistic religions have tended to view these three parts as a hierarchy. God as representative of the divine realm is the most important and most powerful part of reality; his vision and his goodness are responsible for giving order to the human and material realms. Next in order of importance comes the human realm, which has certain things in common with the divine realm. Last in order is the material realm, which should be subdued or stewarded. Evil arises in part when the material realm gains ascendancy over human beings (when it is idolized). Although in this view the divine and instinctive parts of the mind are seen as opposing one another, it is also true that they have something important in common. Of the three parts of the mind, only one part is distinctly human and distinctly under the control of human beings: what I call “reflection.” Alternatively, both the divine and instinctive parts are in some sense “given” to human beings in the form of sensory and non-sensory perceptions, intuitions, observed patterns, external forces etc. These are all decidedly alien, outside, other.

In modern philosophy this tripartite view of the human mind (and of the world) is largely retained, but the emphasis is reversed. Nietzsche and Freud were both primarily interested in the power of instinct. They saw that religious ideas and sentiments often arise from the repression of instinct. The truth and importance of this observation is attested by its continued vitality, and Jung agreed with Freud that often “unconscious” instincts and drives are repressed and find sublimation in religion and other things. But he also saw in his patients that just as often instincts and “biological” drives gain their power from repressed morality and spirituality.^{5} This latter insight seems to be of some relevance with regard to understanding recent history and the psychiatric condition of modern people. Jung’s basic psychiatric position could be formulated as follows: there is no single drive that is fundamental, to which all other drives can be reduced. Instead each drive is a manifestation of “libido” (i.e. psychic energy^{6}) finding the easiest and most appropriate outlet in a person in a particular situation. Here we have one of the first formulations of what could be called a nonlinear theory of psychology. The meaning of a particular neurosis, for example, cannot be pinned down to a particular concrete drive or wish or deficiency; rather it is the manifestation of the Unconscious as a whole – a response to the total situation of the individual, and as such it may have many simultaneous functions and meanings.

Because of this more generalized theory of libido, Jung began to see that the old view of the hierarchy of the mind and world might in fact conceal a hidden unity. The instinctive and the spiritual might have more than an accidental resemblance to one another; instead they might actually be the same thing looked at in different ways. For Jung the tripartite hierarchy of the human mind became circular. The point at which instinct is at its very deepest, and the point at which spirit is at its highest, are both the most radically inhuman. These two for Jung represent the same point, which lies on a circle opposite from the purely human. This way of looking at things has the potential to lead to a very wide and unified theory of psychology and religion throughout history.

5Many of Jung’s case studies involve neuroses that are resolved through some form of mythical or religious development, e.g. “Two Essays on Analytical Psychology.” in Collected Works Vol.7, 81-88, 133. For a discussion of how destructiveness results from the repression of religiosity, see “On the Psychology of the Unconscious,” in Collected Works Vol.7, 94. Also, “Concerning the Archetypes and the Anima Concept,” in Collected Works Vol.9 Part 1, 62-3.

6For example, see C.G. Jung “On Psychic Energy,” in Collected Works Vol. 8.

Anderson 6

In this particular essay I have intended to examine, in accordance with its principle question, how much of Jung’s view of the Unconscious we can make sense of in the context of orthodox evolutionary theory. By putting emphasis on that particular creation myth, we sacrifice some of the potential generality of Jung’s theory. At the same time, the thought experiment provides an interesting possible corrective to the negative view taken by many people of the relationship between human beings and their evolutionary past.

What makes the spiritual and instinctive parts of the mind seem so different in the traditional view is that the spiritual seems to be filled with a sense of wisdom and meaningfulness, while the instincts, though useful, are often blind and fallible. Orthodox evolutionary theory has exaggerated this apparent difference by emphasizing the lack of aim or purpose behind natural selection (which created the instincts and the material world). In the pseudo-Jungian view that I am proposing here, however, the lack of aim apparent in the creation of instincts must also be applied to the creation of the spiritual side of human beings. Evolutionary theory does not simply refute God as the creator of the universe – thereby taking away all spiritual meaning in the world. Rather, evolutionary theory is the creative power that lies behind even the spiritual side of reality.^{7} By looking at religious experience this way, we are in effect moving God inside the psyche of humans and other sentient beings – to the same somewhat nebulous place where instincts are usually envisioned. Within this created nature, the three parts of the human psyche still retain many of the properties that were traditionally ascribed to them, but there are also some differences. For one thing the Divine is no longer unambiguously omnipotent and omniscient, at least not in any complete and perfect way.

Before Jung’s theory, this arrangement was less tenable. It seemed as though a design process like evolution was not sophisticated enough to create God or a godlike power within the psyche, unless God was viewed merely as a kind of delusion or wish fulfillment. In the view I am proposing here, however, evolutionary theory has actually provided us with an excellent model, possibly for the first time in history, of just the kind of alien wisdom that is so characteristic of God and the divine realm. To explain, let me attempt a description of Jung’s view of the simultaneously wise and blind working of the Unconscious, reflecting the hypothesis that the Unconscious was created by a nonlinear design process.

The mind, for Jung, is not a blank slate; rather our subjective experiences . our sensations, desires, thoughts, etc . are the creations of complex psychic (or neurological) structures.^{8} A whole world of conscious experience exists within the mind of each human being, and nothing we do can pierce through the boundary of the individual psyche. To say so is not to belittle this world, or to discredit it as mere fantasy. Human experience just doesn’t make sense without this qualification. At the same time, as soon as we take this view of psychic life, we are thrown back upon the question of where our experiences come from and what lies beneath or within the world as we know it. Since every experience that a person has occurs in the present, and happens within the psyche, it is impossible definitively to separate out the objective historical past from the individual past. From a purely logical point of view, our empirical observations are no more likely to reveal the objective history of the universe than they are to reveal the psychic history of the individual.

7This seems to be overstepping the more cautious methods utilized by biologically oriented psychologists and philosophers such as Daniel Dennett. Dennett’s work has been extremely helpful in its clarity of argument and the timeliness of its message to re-evaluate the foundations of philosophy in the light of the theory of evolution. However, I think that it is still very fruitful to take an “auto-phenomenological” approach as well.

8For example, see C.G. Jung, Memories, Dreams, Reflections, 348. Also “The Archetypes and the Collective Unconscious,” in Collected Works Vol. 9 Part 1, 66, 77.

Anderson 7

These two histories are distinct, and yet they both must be inferred from data gathered in the same field. For Jung, the disharmony of these two competing views, subjective and objective, is interpreted not as evidence of psychological error, but rather as psychological fact.^{9}

The psychic entities that create and shape the world of experience are referred to by Jung as “archetypes.” These archetypes stand somewhere in between subjective and objective experience. In the tradition of analytical psychology the archetypes are frequently given evocative names such as Mother, Father, Child, Hero, Soul, Sun, God, etc; and anecdotes are often related in which patients directly experience archetypal images in dreams and visions. But it is misleading to treat archetypes as though they are fixed objects or fixed images. Just as, for example, the brain itself can never be known directly, but only its image as manifested by the brain itself, so too all the archetypes are meant as conceptual place-holders for that underlying reality which, though it cannot be known directly, makes itself known to us indirectly by how it acts upon us. The archetypes are like shadows that the spiritual and instinctive (i.e. objective) parts of the psyche cast upon the reflecting ego.

Although they do not have human personalities, the archetypes relate to the ego in what Jung sees as a kind of psychic conversation. Consider, as an analogy, what would happen if a traveler knew only a handful of words in a foreign language, and if a native was trying to convey an important message to him. How much would he be able to understand? It is possible that through an imaginative synthesis of words the native might be able to convey some part of his meaning; or perhaps, though failing to communicate, he could at least rouse the traveler to a desired action, the meaning of which might be completely misunderstood. If the traveler knew more words, then the native might be able to approximate more closely his actual intention. But really it is not so much a matter of knowing more words as it is a matter of knowing the right words. Recalling the Logos of John the Evangelist, it seems as though even a hundred-thousand words might still result in vain approximation, so long as the two interlocutors do not have the right word in common.

Just as the words we speak determine how much we can be told, but not the intention of the message itself, so too the very form of our subjective experience determines how that experience acts on us – whether we receive commands from desire, from revelation, from reason, or not at all. If I believe in one particular kind of desire, but no others, then I am likely to feel that desire all the time. It is likely that the same is true for all the various means by which the Unconscious conveys its messages to us. In the relationship between the ego and the Unconscious, the following formula seems to hold: I do not believe because I experience; I experience because I believe.

All subjective experience therefore, according to this formulation, is archetypal. There may be a qualitative difference between, for example, sensory experience, the feeling of a desire, and a numinous vision in a dream, but all of these experiences are alike in that they are expressions of the Unconscious. In every case the psyche wants us to do something, to feel something, to know something. But it is not possible, in principle, to know what the Unconscious is actually trying to say. I would like to call this the first layer of indeterminacy in subjective experience: we know things in our own terms, and we are given by the Unconscious only what we are capable of receiving. Common sense tells us that our instincts want us to do certain things. We feel hungry because our bodies need food; and of course it is probably true that the psyche presents us with

9The origin of the universe, for example, has this dual nature: it is, on the one hand, an event that seems to have occurred billions of years before the birth of the individual; and, on the other hand, it is an event that seems to have occurred at the exact moment when psychic life arose in the individual . perhaps close to the time of birth. In the former view, the world gives birth to the psyche. In the latter view, the psyche gives birth to the world.

Anderson 8

the feeling of hunger for something like this reason. However, as with the behavior of any system with a nonlinear design, it is unlikely that hunger has a single trigger and a single aim. All we can say, from the point of view of evolutionary design, is that the feeling is a manifestation of the Unconscious as a whole, a response to the total situation of the individual. It is a product of millions of years of evolutionary design, and, as such, we should not be surprised if it has more than one function and meaning. I would like to call this the second layer of indeterminacy of subjective experience: there is not, even in reality, a single interpretation of any experience which captures its whole meaning.^{10}

The reason that some archetypal images, such as those of the Mother and Father, are particularly numinous is not really, as is often said, that these images have been imprinted on the psyche as a result of our continual experience of mothers and fathers throughout evolutionary history. Rather, it seems as though the most numinous experiences are those which capture precisely the multiplicity and ambiguity of the Unconscious, its meanings, motivations, and intentions. Mothers and fathers have always been before our eyes, and so they have always been available as metaphors and symbols for many aspects of psychic life. More generally, all the characteristics of the natural world around us can have, over and above their literal meaning, additional constellations of psychic meanings. ^{11} The same holds for the internal world, including those pure theoretical concepts that each generation must learn or discover for itself.

The wisdom of the Unconscious is much like the wisdom of an ecosystem. A trait of a single organism may seem to have little general importance. But if other traits and other organisms within the ecosystem have adapted to its presence, then the trait might actually be essential to the survival of the whole ecosystem, even if its original function is no longer relevant. Within the scope of the first layer of indeterminacy described above, the particular beliefs and habits of the ego evolve a variety of functions during an individual’s life. Therefore even a small change of mind, a new realization, feeling, or idea, can have serious repercussions within the psychic economy. Within the scope of the second layer of indeterminacy, the Unconscious itself is a web of interrelated actions and reactions, encoded, as it were, within the physical and mental structures of the organism. These structures from our evolutionary past do not lie dormant in forgotten regions of the psyche. They are the forces that are constantly generating the world of subjective experience.

The God archetype, which manifests itself in the concept of the Unconscious itself, seems to be based on the experience of a force or power beyond the natural world as we understand it.^{12} Throughout history people have thought that God is implicitly present whenever something truly miraculous or surprising happens. This led some Enlightenment thinkers to suppose that God depends for his existence upon human ignorance.^{13} He seemed to live only in the unknown. And as knowledge of the world increased, it seemed as though God was forced to flee ever further into

10By “in reality” I mean in the underlying biological reality which I am proposing as a hypothetical foundation of the Unconscious.

11See “Archetypes of the Collective Unconscious” in Collected Works Vol. 9 Part 1, 6 for a discussion of myths concerning natural processes such as sun, moon, summer, winter, etc; these are not “allegories of objective occurrences,”

but rather symbolic expressions of inner facts.

12For a discussion of the God archetype, see “The Relations Between the Ego and the Unconscious,” in Collected Works Vol. 7, 238. Technically Jung identifies the God archetype not with the whole Unconscious but with a specific part of it: “The God-image does not coincide with the unconscious as such, but with a special content of it, namely the archetype of the self. It is this archetype from which we can no longer distinguish the God-image empirically.”

Psychology and Religion: West and East (Collected Works Vol. 11), 468; quoted in Memories, Dreams, Reflections, 395.

13Disraeli expressed the idea succinctly: “Where knowledge ends, religion begins.”

Anderson 9

the dark. By the 20th century our knowledge of the physical universe had grown so much that it was no longer possible for God to exist in outer heaven. He fled to inner-space, to the human psyche, and there he seemed to make his stand upon the last frontier of human ignorance.

The conception of God as a residue of human understanding seems to imply that progress in cognitive science and other empirical disciplines could eventually do away with the experience of God entirely. However, contrary to the outlook of Enlightenment thinkers, there is more to the psyche than theoretical knowledge, and there is more to God than theoretical ignorance.^{14} The Unconscious could be thought of as a kind of separate non-personality that arises from physiological activity within the brain, body, and external world. This non-personality (or maybe personality^{15} ) is psychically disconnected from ego-consciousness, in the sense that under normal circumstances we are not aware that there is any other consciousness present in us but that of the ego. At the same time the Unconscious is materially connected to those parts of the brain that give rise to ego-consciousness. In this sense our “God” is invisible and yet has the power to act upon consciousness, to look in upon the world of conscious experience, and to step out from the veiled divine realm and to act in the “world.”^{16} Precise knowledge of the mechanisms by which the Unconscious operates would not change this fundamental power relationship between the ego and the Unconscious.

Like the traditional divine realm, Jung’s Unconscious can be seen as the source of order and energy in the mind or cosmos. It may not be infallible, omnipotent, and omniscient, like some traditional gods, but it has enough in common with the Divine, including great and inhuman “wisdom,” that we can justly consider it yet another face of God. The Unconscious was an object of devotion for Jung himself, and it was a beacon to him in his psychoanalytic practice. He saw it as the perennial source of will, meaning and morality in people’s lives.

Within the psyche we can hope to find a renewed understanding of religious experience. Yet the spiritual meaning of life can be derived no more from abstract psychological principles than it can be derived from abstract physical and biological principles, or from abstract theological principles. It is something that must be experienced and lived, with fear and doubt. The forces in the depth of the psyche are at times benevolent, but they can also be frightening, cruel, and

14In Psychological Types (Collected Works Vol. 6), 57-64, Jung discusses how psychology as an “intellectual endeavor” limits psychological life.

15The question of whether or not the Unconscious parts of the psyche or brain have their own personalities is difficult to resolve, in particular because of the indeterminacies mention above. Having looked at the Unconscious as a product of an evolutionary design process, it seems as though it might be overly anthropocentric to consider it to have a unified personality with a unified will. But there is a whole spectrum of possible quantities and qualities of personality and will. It is possible, for example, to imagine many layers within the Unconscious, so that the layer which underlies and creates ego-consciousness is a somewhat human-like personality that is itself governed by deeper layers of the Unconscious.

16According to my conception, there seems to be a distinct divinity within each separate individual psyche. This would seem to be a significant departure from the conventional religious opinion that the same divine realm presides over all individuals and the whole objective world. From the point of view of individual subjective experience, one divine realm does indeed preside over all individuals, including the ego-personality. But paradox arises when we try to reconcile this subjective view with the objective evolutionary view that I have developed in this essay. It should be noted that for Jung the “collective unconscious” (which I have not had reason here to distinguish from the “personal unconscious”) is considered identical in every human individual. Cf. “On the Psychology of the Unconscious” in Collected Works Vol. 7, 66: “We have to distinguish between the personal unconscious and an impersonal or transpersonal conscious. We speak of the latter also as the collective unconscious, because it is detached from anything personal and is common to all men, since its contents can be found everywhere, which is naturally not the case with the personal contents.”

Anderson 10

destructive. In this more than anything else Jung’s psychology resembles religions of the past and present.

Bibliography

James, William, The Varieties of Religious Experience, in Writings: 1902-1910. New York: Literary Classics,

1987. Originally published in 1902. Jung, C.G., Collected Works vol. 6: Psychological Types, A revision / edition. Ed. H. Read et al. Trans.

R.F.C. Hull. Princeton: Princeton University Press, 1976. —, Collected Works vol. 7: Two Essays on Analytical Psychology, 2nd edition. Ed. H. Read et al. Trans.

R.F.C. Hull. Princeton: Princeton University Press, 1966. —, Collected Works vol. 8: The Structure and Dynamics of the Psyche, 2nd edition. Ed. H. Read et al. Trans.

R.F.C. Hull. Princeton: Princeton University Press, 1970. —, Collected Works vol. 9 Part 1: The Archetypes and the Collective Unconscious, 2nd edition. Ed. H. Read

et al. Trans. R.F.C. Hull. Princeton: Princeton University Press, 1968. —, Collected Works vol. 11: Psychology and Religion: West and East, 2nd edition. Ed. H. Read et al. Trans.

R.F.C. Hull. Princeton: Princeton University Press, 1975. —, Memories, Dreams, Reflections, revised edition. Ed. Aniela Jaffe. Trans. Richard and Clara Winston.

New York: Vintage Books, 1989.